OpenStudy (anonymous):

Medal to person who can help me..... Use a graphical method to determine the approximate interval for which the second order Taylor polynomial for ln(1+x) at x=0 approximates ln(1+x) with an absolute error of no more than 0.04... I have no idea where to start. My book does a bad job at explaining series problems--especially this special case.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@dumbcow

OpenStudy (luigi0210):

Sorry, only in calc I.

OpenStudy (anonymous):

Oh well. Thank you very much regardless.

jimthompson5910 (jim_thompson5910):

First you have to find the second order taylor polynomial for ln(1+x) centered at x = 0

jimthompson5910 (jim_thompson5910):

any ideas on how to do that?

OpenStudy (anonymous):

well its derivative right? Well then that entails yo uto find derivative of ln(1+x) which is 1/1+x. Now second order, so find 2nd derivative so now its -(1+x)^-2. Right?

jimthompson5910 (jim_thompson5910):

so far, so good

jimthompson5910 (jim_thompson5910):

now how do you use those derivatives to find the taylor polynomial?

OpenStudy (anonymous):

I'm pretty sure its plug in zero for x since that is what you are subsitituting for. Plus I use McLauren Series cause its centered at x=0..... So the taylor series would be....\[\ln(1+x)=1+-.5x\] Right?

OpenStudy (anonymous):

Just for 2nd order thats why I went only to two right?

jimthompson5910 (jim_thompson5910):

when you plug x = 0 into the first derivative, you get 1/(1+0) = 1/1 = 1 You divide this by 1! = 1 to get 1*1 = 1 Then you multiply that result by (x-0)^1 = x^1 = x So the first term of this taylor polynomial is x

jimthompson5910 (jim_thompson5910):

for the second derivative, you plug in x = 0 to get -1/(1+x)^2 = -1/(1+0)^2 = -1 Then you divide this by 2! = 2*1 = 2 to get 1/2 and then you multiply by (x-0)^2 = x^2 The second term is (-1/2)x^2

jimthompson5910 (jim_thompson5910):

So overall, the 2nd order taylor polynomial is x - (1/2)x^2

jimthompson5910 (jim_thompson5910):

call that T(x), so T(x) = x - (1/2)x^2

OpenStudy (anonymous):

I do not mean to insult you at all in any manner but are you sure?

jimthompson5910 (jim_thompson5910):

yes, I'm sure

jimthompson5910 (jim_thompson5910):

have a look at this page if you're not sure http://tutorial.math.lamar.edu/Classes/CalcII/TaylorSeries.aspx

OpenStudy (anonymous):

ok cool. now what?

jimthompson5910 (jim_thompson5910):

f(x) is the original function, so f(x) = ln(1+x) the difference in the two functions is f(x) - T(x) = ln(1+x) - (x - (1/2)x^2) = ln(1+x) - x + (1/2)x^2

OpenStudy (anonymous):

Why are we subtracting the Taylor polynomial by the function?

jimthompson5910 (jim_thompson5910):

If T(x) perfectly matched up with f(x), then f(x) - T(x) would always be zero but that's not the case. There's error. So that's why f(x) - T(x) is for the most part nonzero

jimthompson5910 (jim_thompson5910):

we subtract to calculate the error E(x) E(x) = |f(x) - T(x)| and we have those absolute value bars in there to make sure the difference is always positive

OpenStudy (anonymous):

oh ok so now the error is ln(1+x) + x + (1/2)x^2 ? Now do I substitute the .04 error in the equation? If so this makes sense haha

jimthompson5910 (jim_thompson5910):

the error is E(x) = |f(x) - T(x)| which is the same as saying E(x) = ln(1+x) + x + (1/2)x^2 we then plug in E(x) = 0.04 since we want the error to be 0.04. Solving for x will give you the values of x that make E(x) = 0.04 true

jimthompson5910 (jim_thompson5910):

sorry I meant to say E(x) = |ln(1+x) + x + (1/2)x^2|

OpenStudy (anonymous):

How do we get the upper and lower limit? The answers given in the multiple choice consists of an inequality involving an upper and lower limit...?

jimthompson5910 (jim_thompson5910):

tell me what you get when you solve |ln(1+x) + x + (1/2)x^2| = 0.04

jimthompson5910 (jim_thompson5910):

you can do so graphically by plotting y1 = |ln(1+x) + x + (1/2)x^2| y2 = 0.04

jimthompson5910 (jim_thompson5910):

or you can plot y1 = |ln(1+x) + x + (1/2)x^2|-0.04 and find the roots

OpenStudy (anonymous):

Oh Ok I see. I understand now. Thank you veyr much. You just solve with one equaling 0.04 and the other -0.04 Am I correct? From there you have the limits. Cool Thank you!!!

jimthompson5910 (jim_thompson5910):

well if you don't use absolute values, then yes, you'll use 0.04 and -0.04

jimthompson5910 (jim_thompson5910):

it will turn out that everything between those two solutions will have the error smaller than 0.04 anything outside that interval will have E(x) > 0.04